# Orbitronics in Two-dimensional Materials

### v1

@@ VII Future Directions
Another promising line of investigation is exploring proximity effects to boost the capabilities of orbitronics. In spintronics, proximity effects have been successfully employed to induce spin-orbit coupling in graphene. Similarly, the close proximity of orbitronic materials to conductive materials can lead to new orbital hybridizations that may facilitate the transport of OAM currents across heterostructures. This approach could open up new possibilities for integrating orbitronics into existing 2D material platforms, enabling efficient OAM transport and expanding the scope of potential applications.
@@
Exploring and optimizing orbital torque in 2D materials is also a promising area for future research. This often involves injecting orbital currents into other materials and therefore, understanding the mechanisms governing the transmission of orbital current across interfaces is of significant interest. The use of magnetic materials, as sources of OAM currents, is also very encouraging. They can generate OAM currents with unconventional polarizations, expanding the possibilities for spin-orbitronic applications [59].

### v2

@@ VII Future Directions
Another promising line of investigation is exploring the manipulation of hybridization between atoms in 2D materials to enhance the capabilities of orbitronics. For instance, pristine graphene is formed by $sp_{2}$-hybridized carbon atoms, resulting in primarily $p_{z}$ states near the Fermi level with no multiorbital character. However, in fully hydrogenated graphene, the carbon atoms undergo $sp_{3}$ hybridization, which causes the valence band to exhibit a multi-orbital character that can be manipulated through strain [23, 171]. Furthermore, proximity effects have been successfully employed in spintronics to induce spin-orbit coupling in graphene [4]. Similarly, the close proximity of orbitronic materials to conductive materials can lead to new orbital hybridizations [111] that may facilitate the transport of OAM currents across heterostructures. This approach could open up new possibilities for integrating orbitronics into existing 2D material platforms, enabling efficient OAM transport and expanding the scope of potential applications.
@@
Exploring and optimizing orbital torque in 2D materials is also a promising area for future research. A key step toward demonstrating orbital torque in 2D systems is to achieve orbital accumulation at the interface with a 2D ferromagnet. This could enable torque via orbital-to-spin conversion, or potentially involve direct coupling to orbital magnetization in systems with appreciable unquenched orbital moments—an interesting direction that remains to be explored. The use of magnetic materials, as sources of OAM currents, is also very encouraging. The breaking of time-reversal symmetry in the Hamiltonian of magnetic materials results in distinct components in the angular momentum conductivity tensor, which would not be present in a nonmagnetic system. These components can be derived using symmetry arguments [59, 172]. They can generate OAM currents with unconventional polarizations, expanding the possibilities for spin-orbitronic applications. A wide range of 2D magnetic materials is currently available, exhibiting electronic properties that range from metallic [173, 174] to insulating [175, 176, 177]. Exploring new phenomena in these materials will undoubtedly be a key focus within the field of orbitronics.
@@
Acknowledgements
The authors thank the Brazilian funding agencies CNPq, CAPES, and FAPERJ for their financial support. TGR acknowledges financial support from FAPERJ, grant numbers E-26/200.959/2022 and E-26/210.100/2023, FCT - Fundação para a Ciência e Tecnologia, project reference numbers UIDB/04650/2020, 2022.06797.PTDC and 2022.07471.CEECIND/ CP1718/CT0001 (with DOI identifier: 10.54499/2022.07471.CEECIND/CP1718/ CT0001) and 2023.11755.PEX (with DOI identifier https://doi.org/10.54499/2023.11755.PEX) . M.C. acknowledges CNPq (Grant No. 317320/2021-1), FAPERJ/Brazil (Grant No. E26/200.240/ 2023) and INCT Materials Informatics for financial support. L.M.C. acknowledges Severo Ochoa Centres of Excellence programme, CEX2021-001214-S and CERCA Programme/ Generalitat de Catalunya. R. B. M acknowledges financial support from INCT of Spintronics and Advanced Magnetic Nanostructures.